OpenStudy (anonymous):

A recipe that calls for 1/2 cups of orange juice and 5/6 cups of lemonade. Nancy modifies the recipe so it can make 8. How many cups of lemonade does Nancy need?

OpenStudy (texaschic101):

How many servings does 1/2 cup of oj and 5/6 cups of lemonade serve originally ?

OpenStudy (anonymous):

I'm not sure

OpenStudy (texaschic101):

Kinda need that to do the problem

OpenStudy (texaschic101):

I wonder if that is just for 1 serving ?

OpenStudy (anonymous):

I think it is I have answer choices

OpenStudy (texaschic101):

IF 1/2 cups oj and 5/6 cup lemonade served 1 person, then... (5/6) / 1 = x/8 --- 5/6 cup lem to 1 person = x cups lem to 8 people cross multiply (1)(x) = (5/6)(8) x = 40/6 x = 6 2/3 cups is 6 2/3 cups of lemonade for 8 people an answer choice ?

OpenStudy (anonymous):

what! that totally made sence and its not A. 5/6 B. 2/3 C. 1 and 1/24 D. 1 and 1/9

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

I am really not sure about this one...I am so sorry

OpenStudy (anonymous):

i can probably help

OpenStudy (anonymous):

yea you are missing info